let alpha and beta be real and z be a complex number. if z^2+alpha Z+ beta=0 has 2 disntinct roots on the line Re(z)=1,then it is necessary that,

Grade:11

1 Answers

Red
5 years ago
Let z1=p+iq and z2=p-iq, as its roots are complex so they are conjugate.Z1+ z2=-alpha2p=-alpha....(1)Real part of z is 1,p=1.Alpha =-2.D1